1) A pistol that fires a signal flare gives it an initial velocity (muzzle
velocity) of 125 m/s at an angle of 55.00(degrees) above the horizontal. You
can ignore air resistance. (Ignore air resistance.)
(a) What is the magnitude of the horizontal and vertical velocity
component of the signal flare?
(b) What is the flare’s maximum height and the distance from its
firing point to its landing point?
(c) How long does the signal flare stays in the air?

Answer by Alan3354(69443)

1) A pistol that fires a signal flare gives it an initial velocity (muzzle
velocity) of 125 m/s at an angle of 55.00(degrees) above the horizontal. You
can ignore air resistance.
-----
(a) What is the magnitude of the horizontal and vertical velocity
component of the signal flare?
---
Horiz= 125*cos(55)
Vert = 125*sin(55)
==============================
(c) How long does the signal flare stays in the air?
----
Using 9.8 m/sec/sec for gravity:
t = 125*sin(55)/9.8 seconds rising, same time falling
---> 125*sin(55)/4.9 seconds =~ 20.897 seconds.
============================
(b) What is the flare’s maximum height
The average vertical speed is 1/2 the vert component of the launch speed,
= 125*sin(55)/2
Max height = (125*sin(55)/2)*20.897 =~ 1069.86 meters
================
and the distance from its firing point to its landing point?
= (flight time)*125*cos(55)